A sales tax on an item is 5%. Express the total cost, C, in terms of the price of the item, P.
Please and thank you!!

Answer by kshlak(4)

P=price of the item
percentage tax on price of item =5%
amount of tax=(5/100)P
total amount=C=price of the item + tax on it
C=P + (5/100)P
C=(105/100)P
C=(21/20)P